find the x-value having the normal distribution N(1000,250) that has 10% of the area under that normal curve lying to its right.
Answer by stanbon(48568)

find the x-value having the normal distribution N(1000,250) that has 10% of the area under that normal curve lying to its right.
------------
1st: Find the z-value that has 10% to the right.
Using a Ti-83 I get InvNorm(0.9) = 1.2816
-----------
2nd: Find the corresponding x-value using z = (x-u)/s
---
1.2816 = (x-1000)/250
x-1000 = 250*1.2816
x-1000 = 320.3879
x = 1320.3879
=====================
Cheers,
Stan H.